# TSQCA: Threshold-Sweep Qualitative Comparative Analysis in R

### v1

@@ 1.3 Contributions of TSQCA
    Multiple solution handling: When Boolean minimization yields multiple logically equivalent solutions, the package automatically identifies Essential Prime Implicants (EPIs) shared across all solutions and Selective Prime Implicants (SPIs) unique to specific solutions.
@@ 3.5 Function overview
otSweep() Outcome threshold sweep sweep_range, thrX
ctSweepS() Single condition threshold sweep sweep_var, sweep_range, thrY
ctSweepM() Multiple condition threshold sweep sweep_list, thrY
dtSweep() Dual threshold sweep sweep_range_Y, sweep_list_X
@@ 5 Demonstration of advanced output and reporting capabilities
This section demonstrates advanced output structures and reporting utilities provided by TSQCA. The purpose is purely technical and pedagogical: to illustrate how TSQCA organizes, summarizes, and reports results obtained from threshold-sweep analyses. All demonstrations use the dataset bundled with the package (sample_data) to ensure full reproducibility and consistency with the CRAN release.
@@ 8.1 Contributions
    Practical tools: Providing four sweep functions (otSweep, ctSweepS, ctSweepM, dtSweep) and automatic report, reducing what would otherwise be a tedious manual process to efficient single-function calls.
@@ 8.3 Future directions
    Graphical visualization: Developing additional visualization functions such as heatmaps for dual sweeps and trajectory plots for tracking solution evolution across thresholds, to complement existing configuration table and report generation capabilities.
@@ 9 Computational details
otSweep() Outcome threshold sweep dat, outcome, conditions, sweep_range, thrX
ctSweepS() Single condition threshold sweep dat, outcome, conditions, thrY, sweep_var, sweep_range
ctSweepM() Multiple condition threshold sweep dat, outcome, conditions, thrY, sweep_list
dtSweep() Dual threshold sweep dat, outcome, conditions, sweep_range_Y, sweep_list_X

### v2

@@ 1.3 Contributions of TSQCA
    Multiple solution handling: For multiple equivalent solutions, the package identifies Essential Prime Implicants (EPIs) and Selective Prime Implicants (SPIs).
@@ 3.1 Design philosophy
3.2 Three types of QCA solutions
QCA minimization can produce three types of solutions depending on how logical remainders (unobserved configurations) are handled. Following standard QCA terminology (Schneider and Wagemann, 2012; Duşa, 2019), TSQCA supports all three solution types through the include and dir.exp parameters:
Table 3: Three types of QCA solutions and their parameter settings
Solution Type include dir.exp Description
Complex (Conservative) "" NULL Most conservative; uses only observed configurations
Parsimonious "?" NULL Most simplified; uses all logical remainders
Intermediate "?" c(1,1,...) Theory-guided; uses remainders consistent with directional expectations
As of version 1.1.0, TSQCA uses the same default parameter values as QCA::minimize(): include = "" and dir.exp = NULL. This produces the complex (conservative) solution by default, ensuring consistency with the established QCA package.
Researchers can obtain other solution types by explicitly specifying these parameters:
# Complex solution (default)
result <- otSweep(dat, outcome, conditions, sweep_range, thrX)
# Parsimonious solution
result <- otSweep(dat, outcome, conditions, sweep_range, thrX,
                  include = "?")
# Intermediate solution
result <- otSweep(dat, outcome, conditions, sweep_range, thrX,
                  include = "?", dir.exp = c(1, 1, 1))
@@ 3.5 S3 methods
      paste(names(x\$settings\$thrX), "=", x\$settings\$thrX,
@@ 3.6 Function overview
otSweep() Outcome threshold sweep sweep_range, thrX, include, dir.exp
ctSweepS() Single condition sweep sweep_var, sweep_range, thrY, dir.exp
ctSweepM() Multi-condition sweep sweep_list, thrY, include, dir.exp
dtSweep() Dual threshold sweep sweep_range_Y, sweep_list_X, dir.exp
@@ 4 Main functions
Note: The examples in this section compute the intermediate solution by explicitly specifying include = "?" and dir.exp = c(1, 1, 1). To obtain the complex solution (default in v1.1.0), omit these parameters.
@@ 5 Demonstration of advanced output and reporting capabilities
This section demonstrates the advanced output structures and reporting utilities of TSQCA. The purpose is technical: to illustrate how TSQCA organizes and reports results from threshold-sweep analyses. All demonstrations use the bundled dataset (sample_data) to ensure reproducibility.
@@ 8.1 Contributions
    Practical tools: Four sweep functions and automatic report generation reduce tedious manual processes to efficient single-function calls.
@@ 8.3 Future directions
    Graphical visualization: Developing visualization functions such as heatmaps and trajectory plots to complement existing output capabilities.
@@ 9 Computational details
otSweep() Outcome threshold sweep dat, outcome, conditions, sweep_range, thrX, dir.exp
ctSweepS() Single condition sweep dat, outcome, conditions, thrY, sweep_var, sweep_range, dir.exp
ctSweepM() Multi-condition sweep dat, outcome, conditions, thrY, sweep_list, dir.exp
dtSweep() Dual threshold sweep dat, outcome, conditions, sweep_range_Y, sweep_list_X, dir.exp